Amy Cola is considering launching a new soft drink product. The beverage will be sold in a variety of different flavors and will be marketed to young adult. In evaluating the proposed project, the company has the following information:
• The project is estimated to last for 4 years and will have a debt ratio for 10%.
• The company will need to purchase new machinery that has an up-front cost of $40 million (incurred at Year 0). The machinery will be fully depreciated on a 4-year straight-line basis with no salvage value.
• To gain more insight, Amy Cola had spent $1 million on marketing research costs.
• Production of the new product will take place in a vacated facility that the company owns. It is currently empty, and Amy Cola does not intend to lease the facility.
• The project will require a $4 million increase in net operating working capital (NOWC) at Year 0. After Year 0, there will be no changes in NOWC, until Year 4 when the project is completed, and the NOWC is fully recovered.
• The company estimates that sales of the new drink will be $27 million each of the next four years.
• Operating costs (excluding depreciation) are expected to be $11 million each year.
• The company’s tax rate is 20%.
• The owners expect a return of 18% per year but the market interest rate charged by commercial banks are 5% per year.
Based on the information above, the company has estimated the project’s free cash flows as shown in the table below and concluded that it should pursue this business opportunity as it is expected to generate a positive net present value (NPV).

Discuss whether the company has conducted the analysis correctly and what the flaws are. Revise the cash flow estimation, re-calculate the net present value (NPV), the internal rate of return (IRR), payback period and provide your recommendations.
Answer #1

Operating cash flow (OCF) each year = income after tax + depreciation

In year 4, the entire working capital investment is recovered.

The amount spent on market research cost is a sunk cost, i.e. it is already incurred in the past and cannot be recovered. It is not incremental to the acceptance of the project. Hence, it should be ignored in the cash flow analysis.

Cost of capital = (weight of debt * after tax cost of debt) + (weight of equity * cost of equity).

after tax cost of debt = interest rate * (1 - tax rate) = 5% * (1 - 20%) = 4%.

Cost of capital = (10% * 4%) + (90% * 18%) = 16.6%

NPV and IRR are calculated using NPV and IRR functions in Excel

Payback period is the time taken for the cumulative cash flows to equal zero

Payback period = 2 + (cash flow required in year for cumulative cash flows to equal zero / year 3 cash flow) = 2 + ( $14,400,000 / $14,800,000) = 2.97 years.

NPV is -$914,014.

IRR is 15.58%.

It should not accept this project as the NPV is negative and the IRR is lower than the cost of capital.